# On robustness of Spectral Rényi divergence

### v3

@@ 3 Optimization paths
This section presents a further robustness property of the minimum spectral Rényi divergence estimator. We show that (1) the optimization for the minimum spectral Rényi divergence estimator is robust against to outliers in the frequency domain, whereas (2) the optimization for the minimum Itakura–Saito divergence estimator is sensitive to them.
@@ 3.1 Stable optimization path of the Spectral Rényi divergence minimization
We begin with showing that the optimization for the minimum spectral Rényi divergence estimator is robust against to outliers in the frequency domain. Consider the gradient descent update with a sequence of learning rates $\gamma_{1},\gamma_{2},\ldots>0$ of $\widehat{\theta}_{\alpha}$: for the $(m+1)$-th step,
@@
For theoretical results, we make the following assumptions.
@@
We assume the following:
  1. (1)
    The spectral density $S_{\theta}(\omega)$ is differentiable with respect to $\theta$ and we have
    $\sup_{\theta\in\Theta,\,\omega\in[-\pi,\pi]}\max\{\|\nabla_{\theta}S_{\theta}(\omega)\|,\,\|\nabla_{\theta}\log S_{\theta}(\omega)\|,\,\|(1/S_{\theta}(\omega))\nabla_{\theta}\log S_{\theta}(\omega)\|\}$
@@
  2. (2)
    The spectral density $S_{\theta}(\omega)$ is twice differentiable with respect to $\theta$ and we have
    $\limsup_{n\to\infty}\left(\sup_{\theta\in\Theta}\frac{1}{n}\sum_{\omega\in\Omega_{n}}\|\nabla^{2}_{\theta}\log S_{\theta}(\omega)\|_{\mathrm{op}}\right)=:U_{2}<\infty,$
    where let $\|\cdot\|_{\mathrm{op}}$ denote the operator norm.
@@
For any $z\geq 0$, $\omega^{*}\in\Omega_{n}$, and $m\in\mathbb{N}$, the $m$-th gradient decent update $\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}]$ lies in $\Theta$.
Under Assumptions 1 and 2, the subsequent theorem indicates that for any value of $\alpha\in(0,1)$, the path of the gradient descent with a prefixed step size for $\widehat{\theta}_{\alpha}$ is robust against the presence of outliers in the frequency domain.
@@
Remark 1 (Constant step size).
Assumption 1 ensures that we can obtain an upper bound of the smoothness constant of $D^{(n)}_{\alpha}[\widetilde{I}_{n}^{z,\omega^{*}}:S_{\theta}]$ with respect to $\theta$ uniformly in $z\geq 0$: for any $z\geq 0$, we have
@@
$\leq\sup_{z\geq 0}\sup_{\theta\in\Theta}\|\nabla_{\theta}^{2}D_{\alpha}^{(n)}[\widetilde{I}_{n}^{z,\omega^{*}}:S_{\theta}]\|_{\mathrm{op}}$
@@
$\left(-\nabla_{\theta}^{2}D^{(n)}_{\alpha}[\widetilde{I}_{n}:S_{\theta]}]\right)$
@@
[…]
@@
If the further assumption called the Kurdyka–Łojasiewicz inequality [18, 16] holds at all stationary points, that is, points satisfying $\nabla_{\theta}D_{\alpha}^{(n)}[\widetilde{I}_{n}^{z,\omega^{*}}:S_{\theta}]=0$ for the function $\theta\mapsto D_{\alpha}^{(n)}[\widetilde{I}_{n}^{z,\omega^{*}}:S_{\theta}]$, the bounded gradient descent update converges to a stationary point; c.f., Theorem $3.2$ of [1]. Theorem 3 implies that the finite update sequences for different $z\geq 0$ become the same and all stationary points for different $z\geq 0$ become the same as $n\to\infty$, which gives an new aspect into the robust spectral analysis and the optimization theory.
Remark 2 (AR(1) model).
Here we check Assumption 1 for a first-order autoregressive model. The model spectral density is
@@
$\theta_{1}=\log\sigma\quad\text{and}\quad\theta_{2}=\log\frac{1+\rho}{1-\rho}$
@@
$\nabla_{\theta}\log S_{\theta}(\omega)=\begin{pmatrix}2\\ (1+\rho)(1-\rho)\{\cos(\omega)-\rho\}/\{1-2\rho\cos(\omega)+\rho^{2}\}\end{pmatrix}.$
@@
for some absolute positive constant $C$. So, Assumption 1 holds for any bounded parameter space $\Theta$ of $\theta$.
Figure 1: Comparison of optimization paths. The optimization paths of the gradient descent for the $AR(1)$ model, with respect to $(\log\sigma,2\{1/(1+\exp(-\rho))-1/2\})$, are presented under the spectral Rényi divergence minimization and the Itakura–Saito divergence minimization. The left figure shows the optimization paths (curves with light colors) of the spectral Rényi divergence minimization that start from the initial value (the gray circle) and terminate at points after $5000$ iterations (the inverted triangles), where the colors indicates different values of the contamination $z$. The right figure shows the corresponding results for the Itakura–Saito divergence minimization.
@@
$X_{t}^{\circ}=X_{t}+\sqrt{z}\sin(t\pi/2),\quad t=1,\ldots,1000$
with $X_{t}$ following from AR(1) model having $(\sigma,\rho)$. We perform 5000 iterations of gradient descent with a step size of 0.01 with respect to
$(\log\sigma,2\{1/(1+\exp(-\rho))-1/2\}).$
The left panel of Figure 1 illustrates the optimization paths and endpoints based on $D_{\alpha=0.5}[I_{1000}:S_{\theta}]$, starting from the initial value (gray circle). Different colors represent different contamination levels $z\in\{0,10^{2},10^{4}\}$. The figure highlights the robustness of the optimization paths under varying contamination levels. In contrast, the right figure of Figure 1 shows the optimization paths and the stopping points based on the Itakura–Saito divergence $D_{1}[I_{1000}\,:\,S_{\theta}]$ starting from the initial value (the gray circle). This figure demonstrates the non-robustness of the optimization paths of the Itakura–Saito divergence minimization under contamination. The instability observed here will be analyzed further in the subsequent subsection.
@@
The proof is given in Appendix C. It employs the descent lemma (c.f., [2]) together with the comparison of gradients.
Remark 3 (Stability of optimization path with the Armijo condition).
Let $\{\gamma_{m}[\widetilde{I}_{n}]:m=1,2,\ldots\}$ and $\{\gamma_{m}[\widetilde{I}^{z,\omega^{*}}_{n}]:m=1,2,\ldots\}$ the step size sequences for $D_{\alpha}^{(n)}[\widetilde{I}_{n}:S_{\theta}]$ and $D_{\alpha}^{(n)}[\widetilde{I}_{n}^{z,\omega^{*}}:S_{\theta}]$, respectively. From Proposition 2, we can assume
@@
3.2 Instable optimization path of the Itakura–Saito divergence minimization
We next show that the optimization for the minimum Itakura–Saito divergence estimator is sensitive to outliers in the frequency domain. Consider the gradient descent update with a sequence of learning rates $\gamma_{1},\gamma_{2},\ldots>0$ of $\widehat{\theta}_{1}$: for the $(m+1)$-th step,
$\widehat{\theta}_{1}^{(m+1)}[\widetilde{I}_{n}]=\widehat{\theta}_{1}^{(m)}[\widetilde{I}_{n}]+\gamma_{m}\,\mathcal{G}_{1}(\widehat{\theta}_{1}^{(m)}[\widetilde{I}_{n}]\,;\,\widetilde{I}_{n})$
@@ 3.2 Instable optimization path of the Itakura–Saito divergence minimization
$\mathcal{G}_{1}(\theta\,;\,\widetilde{I}_{n})=\frac{1}{n}\sum_{\omega\in\Omega_{n}}\left[\frac{\widetilde{I}_{n}(\omega)}{S_{\theta}(\omega)}-1\right]\nabla_{\theta}\log S_{\theta}(\omega).$
[…]
@@
$\inf_{\theta\in\Theta}\|(1/S_{\theta}(\omega^{*}))\nabla_{\theta}\log S_{\theta}(\omega^{*})\|>0$
holds and $\theta^{(0)}$ lies in $\Theta$. Then we have
$\sup_{z>0}\left\|\widehat{\theta}^{(1)}_{1}[\widetilde{I}^{z,\omega^{*}}_{n}]-\widehat{\theta}^{(1)}_{1}[\widetilde{I}_{n}]\right\|=\infty\,\,\text{almost everywhere}.$ (10)
Further, assume that for any $z>0$ and $\omega^{*}\in\Omega_{n}$, the gradient decent sequence $\{\widehat{\theta}^{(m)}_{1}[\widetilde{I}^{z,\omega^{*}}]\}_{m=1,2\ldots}$ lies in $\Theta$. Then, even when the gradient descent sequence with $\widetilde{I}^{z,\omega^{*}}_{n}$ has a subsequence converging to a stationary point $\widehat{\theta}^{(\infty)}_{1}[\widetilde{I}^{z,\omega^{*}}_{n}]$ for $D_{1}^{(n)}[\,\widetilde{I}^{z,\omega^{*}}_{n}\,:\,S_{\theta}\,]$ , i.e.
$\nabla_{\theta=\widehat{\theta}^{(\infty)}_{1}[\widetilde{I}^{z,\omega^{*}}_{n}]}D_{1}^{(n)}[\widetilde{I}_{n}^{z,\omega^{*}}\,:\,S_{\theta}]=0,$
this stationary point $\widehat{\theta}^{(\infty)}_{1}[\widetilde{I}^{z,\omega^{*}}_{n}]$ is not a stationary point for $D_{1}^{(n)}[\,\widetilde{I}_{n}\,:\,S_{\theta}\,]$:
$\sup_{z>0}\left\|\nabla_{\theta=\widehat{\theta}^{(\infty)}[\widetilde{I}_{n}^{z,\omega^{*}}]}D_{1}^{(n)}[\,\widetilde{I}_{n}\,:\,S_{\theta}\,]\right\|=\infty\,\,\text{almost everywhere}.$
@@ 4 Simulation studies
We present numerical studies employing the Brune spectral model with attenuation that is motivated by seismological studies [33]. We compare five estimation methods:
  • the minimum spectral Rényi divergence estimators with

### v4

@@ 3 Optimization paths
This section presents a further robustness property of the minimum spectral Rényi divergence estimator. We show that (1) the optimization for the minimum spectral Rényi divergence estimator is robust against outliers in the frequency domain, whereas (2) the optimization for the minimum Itakura–Saito divergence estimator is sensitive to them.
@@ 3.1 Stable optimization path of the Spectral Rényi divergence minimization
We begin with showing that the optimization for the minimum spectral Rényi divergence estimator is robust against outliers in the frequency domain. Consider the gradient descent update with a sequence of learning rates $\{\gamma_{i}>0:i=1,2,\ldots\}$ of $\widehat{\theta}_{\alpha}$: for the $(m+1)$-th step,
@@
For theoretical results, we make the following assumptions. Let $\|\cdot\|$ denote the Euclidean norm.
@@
Let $S_{\theta}\in\mathcal{S}_{\Theta}$. For each $\omega\in(0,\pi)$, the function $\theta\mapsto S_{\theta}(\omega)$ is twice continuously differentiable on $\mathrm{int}(\Theta)$, where $\mathrm{int}(\Theta)$ denotes the interior of $\Theta$. Moreover, the following conditions hold:
$\sup_{\theta\in\mathrm{int}(\Theta),\,\omega\in[-\pi,\pi]}\max\{\|\nabla_{\theta}S_{\theta}(\omega)\|,\,\|\nabla_{\theta}\log S_{\theta}(\omega)\|,\,\|(1/S_{\theta}(\omega))\nabla_{\theta}\log S_{\theta}(\omega)\|\}$
@@
and
$\limsup_{n\to\infty}\left(\sup_{\theta\in\mathrm{int}(\Theta)}\frac{1}{n}\sum_{\omega\in\Omega_{n}}\|\nabla^{2}_{\theta}\log S_{\theta}(\omega)\|_{\mathrm{op}}\right)=:U_{2}<\infty,$
where $\|\cdot\|_{\mathrm{op}}$ denotes the operator norm.
@@
For any $z\geq 0$, $\omega^{*}\in\Omega_{n}$, and $m\in\mathbb{N}$, the $m$-th gradient descent update $\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}]$ lies in $\mathrm{int}(\Theta)$.
Under Assumptions 1 and 2, the subsequent theorem indicates that for any value of $\alpha\in(0,1)$, the path of the gradient descent with a prespecified step size for $\widehat{\theta}_{\alpha}$ is robust against the presence of outliers in the frequency domain.
@@
Remark 1.
(Constant step size)  Assumption 1 ensures that we can obtain an upper bound of the smoothness constant of $D^{(n)}_{\alpha}[\widetilde{I}_{n}^{z,\omega^{*}}:S_{\theta}]$ with respect to $\theta$ uniformly in $z\geq 0$: for any $z\geq 0$, we have
@@
$\leq\sup_{z\geq 0}\sup_{\theta\in\mathrm{int}(\Theta).}\|\nabla_{\theta}^{2}D_{\alpha}^{(n)}[\widetilde{I}_{n}^{z,\omega^{*}}:S_{\theta}]\|_{\mathrm{op}}$
@@
$\left(-\nabla_{\theta}^{2}D^{(n)}_{\alpha}[\widetilde{I}_{n}:S_{\theta}]\right)$
@@
By the standard theory of the gradient descent algorithm (c.f., Theorem $3.2$ of Nocedal and Wright (2006)), for any $z\geq 0$, the gradient along with the gradient descent update $\{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}]\}_{m=0,1,2,\ldots}$ with a constant step size $\gamma_{m}=L^{-1}$ vanishes:
@@
If the further assumption called the Kurdyka–Łojasiewicz inequality Łojasiewicz (1963); Kurdyka (1998) holds at all stationary points, that is, points satisfying $\nabla_{\theta}D_{\alpha}^{(n)}[\widetilde{I}_{n}^{z,\omega^{*}}:S_{\theta}]=0$ for the function $\theta\mapsto D_{\alpha}^{(n)}[\widetilde{I}_{n}^{z,\omega^{*}}:S_{\theta}]$, the bounded gradient descent update converges to a stationary point; c.f., Theorem $3.2$ of Attouch et al. (2013). Theorem 3 implies that the finite update sequences for different $z\geq 0$ become the same and all stationary points for different $z\geq 0$ become the same as $n\to\infty$, which provides a new perspective on robust spectral analysis and optimization theory.
Remark 2.
(AR(1) model)  Here we check Assumption 1 for a first-order autoregressive model. The model spectral density is
@@
$\theta_{1}=\log\sigma\quad\text{and}\quad\theta_{2}=\,\mathrm{atanh}(\rho)$
with
$\mathrm{atanh}(\rho)=\frac{1}{2}\log\frac{1+\rho}{1-\rho}$
@@
$\nabla_{\theta}\log S_{\theta}(\omega)=\begin{pmatrix}2\\ 2(1-\rho^{2})\{\cos(\omega)-\rho\}/\{1-2\rho\cos(\omega)+\rho^{2}\}\end{pmatrix}$.
@@
for some absolute positive constant $C$. So, Assumption 1 holds on the interior of any compact subset $\Theta\subset(0,\infty)\times(-1,1)$.
Figure 1: Comparison of optimization paths when the length $n$ of a time series is $200$. The optimization paths of the gradient descent for the AR(1) model, with respect to $(\log\sigma,(1-\exp(-2\rho))/(1+\exp(-2\rho)))$, are presented under the spectral Rényi divergence minimization, the Itakura–Saito divergence minimization, and the time-domain $\gamma$-divergence minimization. Panel (a) shows the optimization paths (curves with light colors) of the spectral Rényi divergence minimization that start from the initial value (the gray circle) and terminate at points after $10000$ iterations (the inverted triangles), where the colors indicate different values of the contamination $z$. Panels (b) and (c) show the corresponding results for the Itakura–Saito divergence and time-domain $\gamma$-divergence minimization, respectively.
@@
$X_{t}^{\circ}=X_{t}+\sqrt{z}\sin(t\pi/2),\quad t=1,\ldots,n=200$
with $X_{t}$ generated from AR(1) model having $(\sigma,\rho)$. We perform 10000 iterations of gradient descent with a step size of 0.01 with respect to
$(\log\sigma,(1-\exp(-2\rho))/(1+\exp(-2\rho))).$
Figure 1 (a) illustrates optimization paths and endpoints based on $D^{(n)}_{\alpha=0.5}[I_{n}:S_{\theta}]$, starting from the initial value (gray circle). Different colors represent different contamination levels $z\in\{0,10^{1},10^{2},10^{3}\}$. The figure highlights the robustness of the optimization paths under varying contamination levels. In contrast, Figure 1 (b) shows the optimization paths and the stopping points based on the Itakura–Saito divergence $D^{(n)}_{\mathrm{IS}}[I_{n}\,:\,S_{\theta}]$ starting from the initial value (the gray circle). This figure demonstrates the non-robustness of the optimization paths of the Itakura–Saito divergence minimization under contamination. The instability observed here will be analyzed further in the subsequent subsection. Further, Figure 1 (c) shows the optimization paths and the endpoints based on the time-domain $\gamma$-divergence. Although Theorem 1 establishes a link between the time-domain $\gamma$-divergence and the spectral $\alpha$-Rényi divergence, their non-asymptotic behavior is different and the implementation of the time-domain $\gamma$-divergence minimization becomes increasingly difficult as $n$ grows, because each optimization step requires a costly log-determinant computation; see also the proof of Theorem 1 in the supplement. This comparison highlights a practical advantage of the approach using the spectral Rényi divergence.
@@
The proof is given in Appendix C. It employs the descent lemma (c.f., Bertsekas (1999)) together with the comparison of gradients.
Remark 3.
(Optimization path with the Armijo condition)  Let $\{\gamma_{m}[\widetilde{I}_{n}]:m=1,2,\ldots\}$ and $\{\gamma_{m}[\widetilde{I}^{z,\omega^{*}}_{n}]:m=1,2,\ldots\}$ the step size sequences for $D_{\alpha}^{(n)}[\widetilde{I}_{n}:S_{\theta}]$ and $D_{\alpha}^{(n)}[\widetilde{I}_{n}^{z,\omega^{*}}:S_{\theta}]$, respectively. From Proposition 2, we can assume
@@
3.2 Unstable optimization path of the Itakura–Saito divergence minimization
We next show that the optimization for the minimum Itakura–Saito divergence estimator is sensitive to outliers in the frequency domain. Consider the gradient descent update with a sequence of learning rates $\{\gamma_{i}>0:i=1,2,\ldots\}$ of $\widehat{\theta}_{\mathrm{IS}}$: for the $(m+1)$-th step,
$\widehat{\theta}_{\mathrm{IS}}^{(m+1)}[\widetilde{I}_{n}]=\widehat{\theta}_{\mathrm{IS}}^{(m)}[\widetilde{I}_{n}]+\gamma_{m}\,\mathcal{G}_{\mathrm{IS}}(\widehat{\theta}_{\mathrm{IS}}^{(m)}[\widetilde{I}_{n}]\,;\,\widetilde{I}_{n})$
@@ 3.2 Unstable optimization path of the Itakura–Saito divergence minimization
$\mathcal{G}_{\mathrm{IS}}(\theta\,;\,\widetilde{I}_{n})=\frac{1}{n}\sum_{\omega\in\Omega_{n}}\left[\frac{\widetilde{I}_{n}(\omega)}{S_{\theta}(\omega)}-1\right]\nabla_{\theta}\log S_{\theta}(\omega).$
The subsequent proposition suggests that both the initial update and the convergent point for $\widehat{\theta}_{\mathrm{IS}}$ are highly sensitive to outliers in the frequency domain.
@@
$\inf_{\theta\in\mathrm{int}(\Theta)}\|(1/S_{\theta}(\omega^{*}))\nabla_{\theta}\log S_{\theta}(\omega^{*})\|>0$
holds and $\theta^{(0)}$ lies in $\mathrm{int}(\Theta)$. Then we have
$\sup_{z>0}\left\|\widehat{\theta}^{(1)}_{\mathrm{IS}}[\widetilde{I}^{z,\omega^{*}}_{n}]-\widehat{\theta}^{(1)}_{\mathrm{IS}}[\widetilde{I}_{n}]\right\|=\infty\,\,\text{almost surely}.$ (10)
Further, assume that for any $z>0$ and $\omega^{*}\in\Omega_{n}$, the gradient descent sequence $\{\widehat{\theta}^{(m)}_{\mathrm{IS}}[\widetilde{I}^{z,\omega^{*}}]\}_{m=1,2\ldots}$ lies in $\mathrm{int}(\Theta)$. Then, even when the gradient descent sequence with $\widetilde{I}^{z,\omega^{*}}_{n}$ has a subsequence converging to a stationary point $\widehat{\theta}^{(\infty)}_{\mathrm{IS}}[\widetilde{I}^{z,\omega^{*}}_{n}]$ for $D_{\mathrm{IS}}^{(n)}[\,\widetilde{I}^{z,\omega^{*}}_{n}\,:\,S_{\theta}\,]$, i.e.
$\nabla_{\theta=\widehat{\theta}^{(\infty)}_{\mathrm{IS}}[\widetilde{I}^{z,\omega^{*}}_{n}]}D_{\mathrm{IS}}^{(n)}[\widetilde{I}_{n}^{z,\omega^{*}}\,:\,S_{\theta}]=0,$
this stationary point $\widehat{\theta}^{(\infty)}_{\mathrm{IS}}[\widetilde{I}^{z,\omega^{*}}_{n}]$ is not a stationary point for $D_{\mathrm{IS}}^{(n)}[\,\widetilde{I}_{n}\,:\,S_{\theta}\,]$:
$\sup_{z>0}\left\|\nabla_{\theta=\widehat{\theta}^{(\infty)}[\widetilde{I}_{n}^{z,\omega^{*}}]}D_{\mathrm{IS}}^{(n)}[\,\widetilde{I}_{n}\,:\,S_{\theta}\,]\right\|=\infty\,\,\text{almost surely}.$
@@ 4 Simulation studies
This section presents two numerical studies using the AR(2) model and the Brune spectral model with attenuation that is motivated by seismological studies Yoshimitsu et al. (2023).
4.1 AR(2) model
Table 1: The mean values of estimates with standard deviations for the AR(2) model without any trend. The values closest to the true values are underlined.
$\widehat{\sigma}$ $\widehat{\varphi_{1}}$ $\widehat{\varphi_{2}}$
Rényi with $\alpha=0.50$ 0.83 ($\pm$ 0.06) 0.57 ($\pm$ 0.64) 0.14 ($\pm$ 0.65)
Rényi with $\alpha=0.75$ 0.95 ($\pm$ 0.10) 0.84 ($\pm$ 0.50) -0.17 ($\pm$ 0.44)
Rényi with $\alpha=0.90$ 1.02 ($\pm$ 0.12) 1.01 ($\pm$ 0.40) -0.28 ($\pm$ 0.38)
IS with the periodogram 1.13 ($\pm$ 0.18) 1.09 ($\pm$ 0.48) -0.32 ($\pm$ 0.42)
IS with the smoothed periodogram 1.14 ($\pm$ 0.19) 1.11 ($\pm$ 0.51) -0.34 ($\pm$ 0.44)
Table 2: The mean values of estimates with standard deviations for the AR(2) model with a trigonometric trend. The values closest to the true values are underlined.
$\widehat{\sigma}$ $\widehat{\varphi_{1}}$ $\widehat{\varphi_{2}}$
Rényi with $\alpha=0.50$ 0.85 ($\pm$ 0.05) 0.53 ($\pm$ 0.68) 0.20 ($\pm$ 0.74)
Rényi with $\alpha=0.75$ 0.99 ($\pm$ 0.12) 0.81 ($\pm$ 0.58) -0.14 ($\pm$ 0.51)
Rényi with $\alpha=0.90$ 1.07 ($\pm$ 0.14) 1.02 ($\pm$ 0.46) -0.26 ($\pm$ 0.42)
IS with the periodogram 2.35 ($\pm$ 0.09) 0.02 ($\pm$ 0.05) 0.98 ($\pm$ 0.05)
IS with the smoothed periodogram 2.35 ($\pm$ 0.10) 0.03 ($\pm$ 0.06) 0.97 ($\pm$ 0.06)
First, we check the performance using the AR(2) model, where the spectral density is given by
$S^{\mathrm{AR2}}_{\theta}(\omega):=\sigma^{2}\Big/\left|1-\varphi_{1}e^{-\sqrt{-1}\omega}-\varphi_{2}e^{-2\sqrt{-1}\omega}\right|^{2},\;\;\theta=(\sigma,\varphi_{1},\varphi_{2}).$
We use the true value $\theta^{\ast}:=(\sigma^{\ast},\varphi_{1}^{\ast},\varphi_{2}^{\ast})=(1,0.9,-0.2)$. We use three different initial values: $\theta^{(0),1}=(0.1,0,0)$, $\theta^{(0),2}=(1,0,0)$, and $\theta^{(0),3}=(10,0,0)$. The optimization is constrained to the stationary region of the AR(2) process; see Section E for details. We generate a time series $\mathbf{x}_{n=2^{10}}$ with its spectral density $S_{\theta^{*}}^{\mathrm{AR2}}$ 1000 times, and if we consider the contamination in the frequency domain, we add trigonometric trends to the time series as
$x_{t}^{o}=x_{t}+z_{1}\sqrt{2\pi/n}\sin(A_{1}t)+z_{2}\sqrt{2\pi/n}\sin(A_{2}t),\,\,t=1,\ldots,n$
with $z_{1}=z_{2}=20$, $A_{1}=\pi/4$ and $A_{2}=\pi/8$.
We compare five estimation methods: (1,2,3) the minimum spectral Rényi divergence estimators with